# In-Situ Fine-Tuning of Wildlife Models in IoT-Enabled Camera Traps for Efficient Adaptation

### v2

@@ 4.5. Case Study: Longitudinal Performance
Results on Accuracy. To put the performance of WildFiT into perspective, we also compare it with the following two alternative methods:
@@
Figure 7. The longitudinal system performance on the location ”R09” of D2 dataset.
@@
We observe that WildFiT closely tracks the performance of Oracle, demonstrating the effectiveness of WildFiT’s in-situ fine-tuning in addressing domain shifts. The significant performance gap between Oracle and WildFiT-Syn further supports this point, as WildFiT-Syn does not handle temporal domain shifts. Notably, WildFiT-Syn initially aligns with the performance of WildFiT prior to the first model fine-tuning event triggered by a class distribution shift (indicated by the first green dotted line from the CDD module). After this point, WildFiT-Syn exhibits substantially worse performance over time, highlighting the critical role of continuous adaptation in maintaining model accuracy under real-world domain shifts.
Results on Runtime Performance. We report the end-to-end fine-tuning time for the trace, where a total of 7 fine-tuning events are triggered (two of which are close to each other, forming the second thick red dotted line). These fine-tuning events take 3.7, 7.8, 4.4, 6.9, 3.5, 5.6, and 6.3 minutes, respectively, on the Orin AGX. On the Jetson Xavier NX, they take 16.4, 11.3, 13.6, 14, 11.9, 21.6, and 12 minutes, respectively. For reference, the average time interval between consecutive bursts of frames at this test location is 22.5 minutes, meaning the fine-tuning time remains well below the time gap between frames on both devices.
@@ 4.6. Ablation Studies
NX AGX
FC 80.2 1.27 0.61
FC + Bias 85.3 3.31 1.70
FC + Bias + BatchNorm (ours) 86.9 3.33 1.75
Whole 89.5 4.01 1.99
@@ 5. Related Work
Domain Adaptation. Domain adaptation (DA) focuses on adapting a model trained on the source domain to perform well on a different but related target domain. Typically, DA methods assume the availability of labeled or unlabeled target data for model adaptation, making them a poor fit for wildlife monitoring applications where data collection is labor-intensive. Some DA approaches, known as zero-shot DA, attempt to generalize to unseen target domains by leveraging auxiliary information about the target domain even if they don’t have direct target domain data during training. For example, ZDDA (peng2018zero, ) and CoCoGAN (wang2019conditional, ) learn from the task-irrelevant dual-domain pairs, while Poda (fahes2023poda, ) uses natural language descriptions of the target domain. These approaches cannot be applied, as they rely on auxiliary information that is unavailable for wildlife animal classification tasks.
@@
Data Augmentation. Data augmentation approaches can generally improve the robustness of ML models. The existing literature roughly falls into the following groups: hand-engineered transformations (shi2020towards, ), adversarial attacks (volpi2018generalizing, ), learned augmentation models (huang2017arbitrary, ; choi2019self, ), feature-level augmentation (zhang2017mixup, ; yun2019cutmix, ), and more recently generative models such as diffusion models (trabucco2023effective, ). The most relevant for us are feature-level augmentation techniques such as CutMix (yun2019cutmix, ) and MixUp (zhang2017mixup, ) that generate synthetic images by blending objects of interest with background images. Despite their efficiency, we show that these generic techniques do not produce high-quality synthetic animal images, resulting in limited, if any, performance improvement. Diffusion models, on the other hand, offer the potential for high quality, realistic, image synthesis that can better match target scenes (controlcom, ; objectstitch, ; libcom, ). However, these incur immense computational cost making them infeasible for rapid, in-situ adaptation to changing scenes.
Drift Detection. Drift detection in ML traditionally focuses on identifying changes in data distribution that can impact model performance over time. The literature includes various approaches, such as statistical hypothesis testing, distance-based metrics, and ensemble methods (alibi-detect, ; lipton2018detecting, ). WildFiT leverages statistical methods such as Least-Squares Density Difference (LSDD), which are frequently employed to detect shifts in data streams (rabanser2019failing, ). These methods are chosen for their computational efficiency, making them suitable for execution on IoT devices. What sets WildFiT apart from traditional drift detection is its innovative drift-aware fine-tuning pipeline: monitoring via drift detection, evaluation via performance validation, and then fine-tuning with synthetic data to address domain drifts.
@@ 6. Conclusions
This paper introduces WildFiT, an in-situ fine-tuning system addressing domain shifts for resource-constrained IoT System. The effectiveness of WildFiT in handling both spatial and temporal domain shifts opens up new possibilities for wildlife monitoring in diverse and changing environments. While we focused on wildlife classification in this work, the principles of background-aware data synthesis and drift-aware fine-tuning have broader applicability in other IoT domains where environmental conditions are dynamic and impact how we approach model deployment in resource-constrained, real-world settings.

### v3

@@ 4.5. Case Study: Longitudinal Performance
Results on Accuracy. We compare WildFiT against the following two alternative methods:
@@
Figure 8. Longitudinal system performance on the location ”R09” of D2 dataset. A total of 7 fine-tuning events are triggered as shown by the vertical dotted lines (BDD triggers are in red, and CDD triggers in green). Two BDD events occur close together, forming the second thick red dotted line.
@@
We observe that WildFiT closely tracks the performance of Oracle, demonstrating the effectiveness of WildFiT’s in-situ fine-tuning in addressing domain shifts. The significant performance gap between Oracle and WildFiT-Syn further supports this point, as WildFiT-Syn does not handle temporal domain shifts. Notably, WildFiT-Syn initially aligns with the performance of WildFiT prior to the first model fine-tuning event triggered by a class distribution shift (indicated by the first green dotted line from the CDD module). After this point, WildFiT-Syn exhibits substantially worse performance over time, highlighting the importance of continuous adaptation in dealing with real-world domain shifts.
Fine-Tuning Time. In terms of end-to-end fine-tuning time, a total of 7 fine-tuning events are triggered as shown by the vertical dotted lines in Figure 8. Fine-tuning times range from 2.4 to 5.3 minutes (average: 3.1 minutes) on the Jetson Orin Nano Super, 3.5 to 7.8 minutes (average: 5.4 minutes) on the Orin AGX, and 11.3 to 21.6 minutes (average: 14.3 minutes) on the Jetson Xavier NX. Since the underlying camera trap system is motion-triggered and animals arrive on average every 22.5 minutes at this location, the fine-tuning duration remains well within reasonable bounds relative to the natural arrival rate of wildlife events.
[…]
Components 15W 25W MAXN
Latency Energy Latency Energy Latency Energy
CDD 1.8 ms 15.8 mJ 1.8 ms 22.6 mJ 1.5 ms 21.5 mJ
BDD 68.0 ms 467.2 mJ 70.2 ms 498.3 mJ 63.6 ms 548.6 mJ
Inference 62.8 ms 440.3 mJ 61.1 ms 455.7 mJ 54.9 ms 496.6 mJ
Validation 1.7 s 13.8 J 1.4 s 13.2 J 1.3 s 15.1 J
Fine-Tuning 274.9 s 2634.1 J 191.6 s 2730.4 J 182.3 s 2789.2 J
Table 6. Average latency and energy consumption of each component in the WildFiT pipeline under different power modes, implemented and evaluated on the Jetson Orin Nano (8 GB) for location ”R09” of D2 dataset.
Results on Total Energy Consumption. We now report the total energy consumption of the major components of the WildFiT system, implemented on the Jetson Orin Nano Super (MAXN mode), for location R09 of the D2 dataset, collected via a camera trap over a 37-day period. Figure 9 illustrates the energy breakdown, where approximately one-third to one-fourth of each bar represents system idle energy consumption, with the remainder attributed to active computation. As observed earlier, both the CDD and BDD modules remain highly lightweight in terms of energy usage. Notably, BDD is triggered more frequently due to background dynamics. The total energy consumption for 7 rounds of model fine-tuning was 19.52 kJ. In contrast, the validation module—while incurring energy equivalent to approximately 5 fine-tuning rounds—was triggered nearly 1000 times, successfully filtering out many unnecessary fine-tuning attempts. This highlights its effectiveness in verifying model robustness against minor temporal, spatial, and class distribution shifts, thereby preventing excessive energy expenditure.
Figure 9. Total energy consumption of each component in the end-to-end implementation of WildFiT on the Jetson Orin Nano Super (8 GB) for location R09 of the D2 dataset, collected over a 37-day period. Orange bars indicate the energy consumed by the system while powered on but idle (i.e., not performing any computation).
Results on Battery Requirements. We evaluated energy consumption by running WildFiT on the Jetson Orin Nano Super with a 37-day image stream from the location “R09” in the D2 dataset. The system consumed only 11.2 Wh total, demonstrating exceptional energy efficiency. This low power requirement means the system can operate on just three standard AA batteries, or run for approximately six months on a compact 10,000 mAh power bank, making it highly practical for remote wildlife monitoring deployments.
@@ 4.6. Ablation Studies
Nano NX AGX
FC 80.2 0.79 1.27 0.61
FC + Bias 85.3 1.94 3.31 1.70
FC + Bias + BatchNorm (ours) 86.9 1.98 3.33 1.75
Whole 89.5 2.20 4.01 1.99
@@ 5. Related Work
Domain Adaptation. Domain adaptation (DA) focuses on adapting a model trained on the source domain to perform well on a different but related target domain. Typically, DA methods assume the availability of labeled or unlabeled target data for model adaptation, making them a poor fit for wildlife monitoring applications where data collection is labor-intensive. Some DA approaches, known as zero-shot DA, attempt to generalize to unseen target domains by leveraging auxiliary information about the target domain even if they don’t have direct access to target domain data during training. For example, ZDDA (peng2018zero, ) and CoCoGAN (wang2019conditional, ) learn from the task-irrelevant dual-domain pairs, while Poda (fahes2023poda, ) uses natural language descriptions of the target domain. These approaches cannot be applied, as they rely on auxiliary information that is unavailable for wildlife animal classification tasks. Recently proposed methods, including ZoDi (zodi, ), Min-Max (daynight1, ), and CIConv (daynight2, ), target zero-shot adaptation primarily for day-night or weather-induced shifts. However, these methods assume consistent scene layouts, which are unsuitable for wildlife scenarios and incur heavy computational overhead, making them impractical for resource-constrained deployments.
@@
Data Augmentation. Data augmentation approaches can generally improve the robustness of ML models. The existing literature roughly falls into the following groups: hand-engineered transformations (shi2020towards, ), adversarial attacks (volpi2018generalizing, ), learned augmentation models (huang2017arbitrary, ; choi2019self, ), feature-level augmentation (zhang2017mixup, ; yun2019cutmix, ), and more recently generative models such as diffusion models (trabucco2023effective, ). The most relevant for us are feature-level augmentation techniques such as CutMix (yun2019cutmix, ) and MixUp (zhang2017mixup, ) that generate synthetic images by blending objects of interest with background images. Despite their efficiency, we show that these generic techniques do not produce high-quality synthetic animal images, resulting in limited, if any, performance improvement. Diffusion models, on the other hand, offer the potential for high quality and realistic image synthesis that can better match target scenes (controlcom, ; objectstitch, ; libcom, ; AnyDoor, ). However, these incur immense computational cost making them infeasible for rapid, in-situ adaptation to changing scenes.
Drift Detection. Drift detection in ML traditionally focuses on identifying changes in data distribution that can impact model performance over time. The literature includes various approaches, such as statistical hypothesis testing, distance-based metrics, and ensemble methods (alibi-detect, ; lipton2018detecting, ). WildFiT leverages statistical methods such as Least-Squares Density Difference (LSDD), which are frequently employed to detect shifts in data streams (rabanser2019failing, ). These methods are chosen for their computational efficiency, making them suitable for execution on IoT devices. What sets WildFiT apart is its complete pipeline that combines drift detection, validation, and synthetic data fine-tuning to autonomously address domain shifts.
@@ 6. Conclusions
This paper introduces WildFiT, an in-situ fine-tuning system addressing domain shifts for resource-constrained IoT systems. The effectiveness of WildFiT in handling both spatial and temporal domain shifts opens up new possibilities for wildlife monitoring in diverse and changing environments. While we focused on wildlife classification in this work, the principles of background-aware data synthesis and drift-aware fine-tuning have broader applicability in other IoT domains where environmental conditions are dynamic and impact how we approach model deployment in resource-constrained, real-world settings.